A 7% semiannual coupon bond matures in 5 years. The bond has a face value of $1,000 and a current yield of 7.5026%.

What is the bond's price? Do not round intermediate calculations. Round your answer to the nearest cent.
$  

What is the bond's YTM? (Hint: Refer to Footnote 6 for the definition of the current yield and to Table 7.1) Do not round intermediate calculations. Round your answers to two decimal places.
    %
Answer #1

a)

Bond's price=Annual coupon/current yield =70/7.5026% =933.01

b)

Bond's YTM:

=RATE(nper,pmt,pv,fv)

=RATE(5*2,7%/2*1000,-933.01,1000)*2

=8.68%